Solve the IVP [:] =: -:] [:] [0]=[2]
Answer #1

. Soln 2 Let A = -7. 3 -8 The characteristic เd ๆ equation is LA-^ Il = 0 (2-1)(-8-2) - (-21) i-e. = 0 2-X 3 -7 -8-) TT -16 +. R₂ - 3 R 17]*] =10 72 - 772 = 0 ni = 12 :. For = 1 Ky 01 X2 1 -- : for distinct real eigenvalues ac, 2c, 12 and eigen vecto